OpenStudy (anonymous):

The difference between two integers is at least 14. The smaller integer is 2. What is the larger integer?

OpenStudy (texaschic101):

a = smaller number and b = larger number a = 2 a - b >= 14 2 - b >= 14 -b >= 14 - 2 -b >= 12 b <= 12 you might need a second opinion on this